AFN equation

Broussard Skateboard's sales are expected to increase by 25% from $9.0 million in 2016 to $11.25 million in 2017. Its assets totaled $5 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 4%, and the forecasted payout ratio is 65%. What would be the additional funds needed? Do not round intermediate calculations. Round your answer to the nearest dollar.
$_______.

Assume that an otherwise identical firm had $6 million in total assets at the end of 2016. Broussard's capital intensity ratio (A0*/S0) is (HIGHER THAN, LOWER THAN), than the otherwise identical firm; therefore, Broussard is (LESS, MORE, THE SAME), capital intensive - it would require (A SMALLER, A LARGER, THE SAME) increase in total assets to support the increase in sales.
Answer #1

Sales in 2017 = 11,250,000

Profit in 2017 = 4% = 0.4*11,250,000 = 450,000

Dividend payout = 65% = 0.65*450,000 = 292,500

Retained earnings = 450,000=292,500 = 157,500

Increase in assets for 2017 = 5,000,000*0.25 = 1,250,000

Increase in spontaneow liabilties = 450,000*0.25 + 450,000*0.25 = 225,000

Additional funds needed = Increases in assets - Increase inSpontaneous liabilities - addition to retained earnings

Additional funds needed = 1,250,000 -225000-157500 = 867,500

Addtional funds needed = $867,500

Broussard's capital intensity ratio (A0*/S0) is LOWER THAN the otherwise identical firm; therefore, Broussard is LESS capital intensive - it would require A SMALLER increase in total assets to support the increase in sales.